OpenStudy (anonymous):

Please help ? Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 10 3x + 4y = 25

OpenStudy (anonymous):

-4(2x+y)=-4*10 -8x-4y=-40 3x+4y=25 (-8x-4y)+(3x+4y)=-40+25 -8x-4y+3x+4y=-15 -5x=-15 x=3 3*3+4y=25 9+4y=25 4y=16 y=4 i wanna have sex with you :)

OpenStudy (anonymous):

lmao well what would the real world scenario be then ?

OpenStudy (anonymous):

and maybe you also want :)

OpenStudy (anonymous):

i need a real world scenario for now

OpenStudy (anonymous):

it's a real world scenario we just need to meet :)